OpenStudy (anonymous):

1\2 plus 2\9 what is the sum?

OpenStudy (anonymous):

1/2=9/18 and 2/9=4/18 So when you say 1/2 + 2/9 it is same as 9/18 + 4/18 can you add 9/18 and 4/18 ?

OpenStudy (anonymous):

Its 1\2 plus 4\9 typed it in wrong sorry

OpenStudy (anonymous):

you get the idea, find the common denominator for both fractions 1/2=9/18 and 4/9=8/18 now add the fractions.

OpenStudy (anonymous):

17\18

OpenStudy (anonymous):

YEs, there you go

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

ANYTIME !

OpenStudy (anonymous):

:)